If the probability for passing a test (p) is 85% and variance is 0,638, what is the chance that 3 students in a group will not pass the test?
Answer by stanbon(75887)

If the probability for passing a test (p) is 85% and variance is 0.638, what is the chance that 3 students in a group will not pass the test?
------
Variance = p*q*n
0.638 = 0.85*0.15*n
n = 5
-------
P(x = 3) = 5C3*0.15^3*0.85^2 = binompdf(5,0.15,3) = 0.0244
------------
Cheers,
Stan H.
---------------